### v1

# Rethinking Inter-LoRA Orthogonality in Adapter Merging: Insights from Orthogonal Monte Carlo Dropout

Abstract
We propose Orthogonal Monte Carlo Dropout, a mechanism that enforces strict orthogonality when combining sparse semantic vectors without extra time complexity. LoRA, a popular fine-tuning method for large models, typically trains a module to represent a specific concept such as an object or a style. When multiple LoRAs are merged, for example to generate an object in a particular style, their semantic vectors may interfere with each other. Our method guarantees, at the theoretical and runtime levels, that merged LoRAs remain orthogonal and thus free from direct interference. However, empirical analysis reveals that such orthogonality does not lead to the semantic disentanglement or compositionality highlighted in prior work on compositional adaptation. This finding suggests that inter-LoRA orthogonality alone may be insufficient for achieving true semantic compositionality, prompting a re-examination of its role in adapter merging. Code can be found at https://github.com/andiac/Orth_MC_Dropout
Index Terms—  Low-Rank Adaptation, Monte Carlo Dropout, Diffusion Models, Orthogonality
1 Introduction
Low-Rank Adaptation (LoRA) [1] has become a popular method for adapting large-scale models efficiently by injecting trainable low-rank matrices while keeping pre-trained weights frozen. Beyond single-task adaptation, practitioners often merge multiple LoRAs [2, 3, 4, 5, 6] — e.g., combining an object identity with a visual style — for compositional generation. However, prior studies [7] assume that such merging may suffer from interference when LoRAs are highly aligned, limiting semantic controllability.
Existing approaches such as Zip-LoRA [7] attempt to address this issue by penalizing similarity between LoRA weights, but they neither enforce strict orthogonality nor avoid the need for additional fine-tuning. To overcome these limitations, we propose Orthogonal Monte Carlo (MC) Dropout, a simple yet theoretically grounded mechanism that enforces orthogonality among LoRAs without additional computational overhead. Our analysis proves that the merged updates remain orthogonal at runtime, effectively eliminating direct interference.
Interestingly, experiments reveal that strict orthogonality alone does not yield the semantic disentanglement assumed in earlier compositional adaptation work [7]. This finding highlights a gap between structural orthogonality and true semantic compositionality, motivating a rethinking of the role of inter-LoRA orthogonality in adapter merging.
Our main contributions are as follows:
  • We propose Orthogonal Monte Carlo Dropout, a novel mechanism that enforces strict orthogonality when merging LoRAs, without incurring extra computational cost.
  • We provide theoretical guarantees of runtime orthogonality.
  • Through empirical analysis, we demonstrate that orthogonality does not necessarily imply semantic disentanglement or compositionality, challenging a key assumption in prior compositional adaptation work and prompting a rethinking of inter-LoRA orthogonality in adapter merging.
2 Preliminary
2.1 Monte-Carlo Dropout
Monte-Carlo (MC) Dropout [8] leverages dropout [9], a regularization method that randomly drops units in a neural network. In MC Dropout, dropout is applied during both training and inference stages. By performing multiple forward passes through the network with dropout enabled, it generates a distribution of outputs. This distribution allows for the calculation of predictive uncertainty, providing insights into the confidence of the model’s predictions.
Let $W\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ be the parameters of a linear layer, $p$ be the dropout rate, and $h\in\mathbb{R}^{d_{\text{in}}}$ be the input to the layer. We define a dropout mask $z=[z_{1},\dots,z_{d_{\text{out}}}]^{T}$, where each $z_{i}\sim\text{Ber}(1-p)$ for $i\in\{1,\dots,d_{\text{out}}\}$. Then, the feed-forward function $f(h)$ can be expressed as:
$f(h)=\frac{(Wh)\odot z}{1-p},$ (1)
where $\odot$ denotes element-wise multiplication. Note that the denominator 1p1-p rescales the output to compensate for the expected reduction in magnitude caused by dropout. In this work, we utilize MC Dropout’s property of preserving semantic representation even when (at least) more than half of neurons are masked out [10, 11], rather than exploiting its uncertainty estimation capabilities.
[…]
Fig. 1: Comparison of LoRA merging strategies. In this example, $w_{i}=1$ and $p_{i}=1/3$ for $i\in\{1,2,3\}$. (Top Left) Direct merge simply sums LoRA updates, which may cause interference when updates are aligned. (Top Right) Monte Carlo Dropout merge. (Bottom) Our proposed Orthogonal Monte Carlo Dropout guarantees orthogonality among LoRA contributions, ensuring no direct overlap.
2.2 Low-Rank Adaptation (LoRA)
Low-Rank Adaptation [1] (LoRA) is a technique that involves freezing the pre-trained model weights and injecting trainable rank decomposition matrices into each layer of the attention mechanism. This approach significantly reduces the number of trainable parameters required for downstream tasks.
Let $W$ represent the pre-trained parameters of a linear layer within an attention mechanism, and let $h$ be the input to this linear layer. A pair of LoRA parameters, $\Delta W_{a}$ and $\Delta W_{b}$ (the trainable rank decomposition matrices), operate as follows:
$f(h)=Wh+\Delta W_{b}\Delta W_{a}h.$
Since the multiplication of $\Delta W_{a}$ and $\Delta W_{b}$ forms a low-rank matrix, we define $\Delta W=\Delta W_{b}\Delta W_{a}$. Thus, the operation simplifies to:
$f(h)=Wh+\Delta Wh.$
During the training stage, $W$ is kept fixed while $\Delta W$ is tuned to maximize the likelihood on the fine-tuning dataset.
2.3 Merging LoRAs
A common method to merge two LoRAs, $\Delta W_{1}$ and $\Delta W_{2}$, is through a weighted sum:
$f(h)=Wh+w_{1}\Delta W_{1}h+w_{2}\Delta W_{2}h,$ (2)
where $w_{1}$ and $w_{2}$ are the respective weights. The work of [7] indicates that highly aligned LoRA weights merge poorly. This can be understood straightforwardly: if the vectors $\Delta W_{1}h$ and $\Delta W_{2}h$ point in similar directions, the addition operation will cause them to interfere with each other, ultimately diminishing the effectiveness of the sum.
3 Orthogonal Monte-Carlo Dropout
We propose Orthogonal Monte-Carlo (MC) Dropout to mitigate interference between LoRAs. Let $W$ be the pre-trained parameters, and $\Delta W_{1}$ and $\Delta W_{2}$ be two LoRAs with corresponding dropout rates $p_{1}$ and $p_{2}$, where $1-p_{1}+1-p_{2}\leq 1$. The feed-forward function $f(h)$ is defined as:
$f(h)=Wh+\frac{w_{1}z^{(1)}\odot(\Delta W_{1}h)}{1-p_{1}}+\frac{w_{2}(\mathbf{1}-z^{(1)})\odot z^{(2)}\odot(\Delta W_{2}h)}{1-p_{2}},$
where $z^{(1)}_{i}{\sim}Ber(1-p_{1})$, $z^{(2)}_{i}{\sim}Ber((1-p_{2})/p_{1})$ for $i\in\{1,\dots,d_{\text{out}}\}$. By construction, interference between the two LoRAs is eliminated, since the inner product of $z^{(1)}\odot(\Delta W_{1}h)$ and $(\mathbf{1}-z^{(1)})\odot z^{(2)}\odot(\Delta W_{2}h)$ is zero. For consistency, note that $(1-z^{(1)}_{i})\cdot z^{(2)}_{i}\sim\text{Ber}(1-p_{2})$.
This concept can be expanded to $p_{1},\dots,p_{k}$ for merging $k$ LoRAs, where $\sum_{i=j}^{k}(1-p_{j})\leq 1$. For $k$ LoRAs, the feed-forward function $f(h)$ is defined as:
$f(h)=Wh+\sum_{j=1}^{k}\frac{\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)}\odot(\Delta W_{j}h)}{1-p_{j}},$
where $z^{(1)}_{i}\sim\text{Ber}(1-p_{1})$ and each
$z^{(j)}_{i}\sim\text{Ber}\left((1-p_{j})/\left(\left(\sum_{l=1}^{j-1}p_{l}\right)-(j-2)\right)\right)$
for $i\in\{1,\dots,d_{\text{out}}\}$, $j\in\{2,\dots,k\}$. In the remainder of this section, we demonstrate the consistency and orthogonality of this construction.
Theorem 1.
(Consistency) Let the random mask
$m^{(j)}=\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)},$
then $m^{(j)}_{i}\sim\text{Ber}(1-p_{j})$.
Proof.
Let $i\in\{1,\dots,d_{\text{out}}\}$, $j\in\{1,\dots,k\}$,
$\mathbb{P}(m^{(j)}_{i}=1)$
$=$ $\mathbb{P}(\left(\prod_{l=1}^{j-1}(1-z^{(l)}_{i})\right)\cdot z^{(j)}_{i}=1)$
$=$ $\mathbb{P}(z^{(1)}_{i}=0,\dots z^{(j-1)}_{i}=0,z^{(j)}_{i}=1)$
$=$ $\mathbb{P}(z^{(1)}_{i}=0)\cdots\mathbb{P}(z^{(j-1)}_{i}=0)\mathbb{P}(z^{(j)}_{i}=1)$
$=$ $p_{1}\left(1-\frac{1-p_{2}}{p_{1}}\right)\left(1-\frac{1-p_{3}}{p_{1}+p_{2}-1}\right)\cdots$
$\left(1-\frac{1-p_{j-1}}{\left(\sum_{l=1}^{j-2}p_{l}\right)-(j-3)}\right)\frac{1-p_{j}}{\left(\sum_{l=1}^{j-1}p_{l}\right)-(j-2)}$
$=$ $p_{1}\frac{p_{1}+p_{2}-1}{p_{1}}\cdot\frac{p_{1}+p_{2}+p_{3}-2}{p_{1}+p_{2}-1}\cdots$
$\frac{\left(\sum_{l=1}^{j-1}p_{l}\right)-(j-2)}{\left(\sum_{l=1}^{j-2}p_{l}\right)-(j-3)}\cdot\frac{1-p_{j}}{\left(\sum_{l=1}^{j-1}p_{l}\right)-(j-2)}$
$=$ $1-p_{j}$
To facilitate the proof of orthogonality, we begin with the following proposition:
Proposition 1.
Let $m,n\in\{0,1\}^{d}$ for some dimension $d$, such that $m^{T}n=0$. Then for any $a,b\in\mathbb{R}^{d}$, we have $(m\odot a)^{T}(n\odot b)=0$.
Proof.
As $m$ and $n$ only take values 0 or 1, $m_{i}n_{i}=0$ or $1$ for all $i$. Given $m^{T}n=m_{1}n_{1}+\dots+m_{d}n_{d}=0$, we must have $m_{i}n_{i}=0$ for all $i$ (if any $m_{i}n_{i}=1$, then $m^{T}n>0$). Therefore:
$(m\odot a)^{T}(n\odot b)=m_{1}a_{1}n_{1}b_{1}+\dots m_{d}a_{d}n_{d}b_{d}=0+\dots+0=0$
Lastly, we prove orthogonality:
Theorem 2.
(Orthogonality) For any $i,j\in\{1,\dots,k\}$ with $i<j$, the vectors
$m^{(j)}\odot(\Delta W_{j}h)=\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)}\odot(\Delta W_{j}h)$
and
$m^{(i)}\odot(\Delta W_{i}h)=\left(\prod_{l=1}^{i-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(i)}\odot(\Delta W_{i}h)$
are orthogonal, where $h\in\mathbb{R}^{d_{\text{in}}}$ is an input vector and $\odot$ denotes element-wise multiplication.
Proof.
Each element of $z^{(i)}$ is sampled from a Bernoulli distribution, taking only values 0 or 1. Consequently, for all $j$:
$(1-z_{j}^{(i)})\cdot z_{j}^{(i)}=0$
Therefore:
$(\mathbf{1}-z^{(i)})^{T}z^{(i)}=0$
Then, the inner product of the two vectors
$(m^{(j)}\odot(\Delta W_{j}h))^{T}(m^{(i)}\odot(\Delta W_{i}h))$
$=$ $\left(\left(\prod_{l=1}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)}\odot(\Delta W_{j}h)\right)^{T}$
$\left(\left(\prod_{l=1}^{i-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(i)}\odot(\Delta W_{i}h)\right)$
$=$ $\left((\mathbf{1}-z^{(i)})\odot\left(\prod_{l=1,l\neq i}^{j-1}(\mathbf{1}-z^{(l)})\right)\odot z^{(j)}\odot(\Delta W_{j}h)\right)^{T}$
$\left(z^{(i)}\odot\left(\prod_{l=1}^{i-1}(\mathbf{1}-z^{(l)})\right)\odot(\Delta W_{i}h)\right)$
$=$ $0$
by proposition 1. ∎
4 Experiments
In our experiments, we compare three LoRA merging methods: Direct Merge (introduced in Equation (2)), Monte Carlo Dropout (introduced in Equation (1)), and our proposed Orthogonal Monte Carlo Dropout. A detailed illustration of the differences among these methods is provided in Fig. 1. We include Monte Carlo Dropout as a baseline to control for the effect of dropout itself, thereby isolating the impact of orthogonality. Without this comparison, it would be unclear whether the differences in generated images arise from dropout or from orthogonality.
4.1 Image Generation Settings
Unless otherwise specified, we use SDXL 1.0 [12] as the base model for image generation, producing images at a resolution of $\times1024$ with 40 denoising steps, a clip skip of 2, and a classifier-free guidance (CFG) scale of 7 [13]. Since diffusion-based image generation is highly sensitive to the initial noise [14], we fix the random seed for the initial noise across comparisons to rigorously isolate the effects of orthogonality.
4.2 Merging LoRAs Trained on Dreambooth Datasets
We use the 30 concepts provided in the DreamBooth dataset [15], training a separate LoRA for each concept. Each LoRA is trained with a rank of 128 using the AdamW optimizer, with a learning rate of $10^{-4}$ for the UNet and $10^{-5}$ for the text encoder. We then merge them pairwise using the three methods described above and evaluate the generated images. For each pairwise combination, we set $w_{1}=w_{2}=1$ and $p_{1}=p_{2}=0.5$, and generate four images, yielding a total of $(30\times 29)/2\times 4=1740$ images. As shown in Fig. 2, clear semantic differences among the three methods are difficult to discern by visual inspection, though we observe a noticeable decline in image quality for both Monte Carlo Dropout and Orthogonal Monte Carlo Dropout. To quantify our observations, we use the CLIP score [16] to measure semantic similarity, conduct an A/B test user study to compare the fidelity of concept representation, and further assess the quality of the generated images.
In the user study, we evaluated a total of $1740\times 3=5220$ cases. Each case presented the reference images of two concepts, followed by two generated images from the compared methods. Participants were asked to select which image better represented both concepts. Following [17, 18], we recruited five participants and determined the outcome for each case by majority vote. When both methods received approximately 50% of the votes, we considered the difference indistinguishable to participants.
As shown in Table 1 and Table 2, all three merging strategies achieve similarly high semantic alignment (CLIP >> 0.91), while both MC Dropout and Orthogonal MC Dropout consistently reduce image quality, consistent with our earlier observations. The user study results in Table 1 further indicate that participants did not perceive a difference between Dropout and Orthogonal, but expressed a preference for Direct Merge, which may be attributable to its higher image quality rather than semantic factors.
$\Delta W_{1}$$\Delta W_{2}$DirectDropoutOrthogonalRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Fig. 2: Examples of DreamBooth LoRA merges. ‘Direct’ denotes Direct Merge, ‘Dropout’ denotes Monte Carlo Dropout, and ‘Orthogonal’ denotes Orthogonal Monte Carlo Dropout.
$\Delta W_{1}$$\Delta W_{2}$DirectDropoutOrthogonalRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Fig. 3: Merging LoRAs from the community.
Direct vs Dropout Direct vs Orthogonal Dropout vs Orthogonal CLIP [16] 0.9116 0.9154 0.9346 User Study 84.0% 81.3% 47.3%
Table 1: Quantatitive result on similarity and fidelity.
4.3 Merging LoRAs from Community Sources
In addition to the controlled quantitative experiments presented in the previous section, Fig. 3 shows results from merging LoRAs downloaded from the community. Since our merging operates in the hidden space, LoRAs with different ranks can also be freely combined. These results demonstrate that our proposed method requires no additional fine-tuning and generalizes well, while further highlighting that orthogonality plays a limited role in LoRA merging.
5 Discussions
5.1 Interference as a Source of Image Quality Gains
Direct Merge achieves substantially higher image quality than the other methods, indicating that the interactions arising from direct addition of LoRAs can, in fact, enhance visual quality. This observation calls into question the general assumption that orthogonality is always more advantageous for merging.
5.2 User Study: Isolating Orthogonality
While participants tended to favor higher-quality images, the Dropout vs Orthogonal case in Table 1 demonstrates that, under comparable image quality, enforcing strict orthogonality does not improve user-perceived concept fidelity.
5.3 Relationship with Zip-LoRA
We do not question the effectiveness of Zip-LoRA [7], but note that its performance likely stems from additional fine-tuning rather than orthogonality. Zip-LoRA does not guarantee orthogonality; it only encourages it through an inner-product penalty in the loss.
Direct Dropout Orthogonal $\uparrow$ MUSIQ [19] 62.8063 59.5810 59.6875 $\uparrow$ ARNIQA [20] 0.7529 0.7398 0.7423 $\uparrow$ NIMA [21] 4.2824 4.1120 4.1148 $\uparrow$ TRes [22] 87.3702 84.0270 84.1839 $\uparrow$ DBCNN [23] 0.6647 0.6419 0.6427
Table 2: Image Quality.
6 Conclusion
We introduced Orthogonal Monte Carlo Dropout to enforce strict orthogonality in LoRA merging. Our experiments show that orthogonality alone offers little benefit, as it neither yields semantic disentanglement nor improves compositionality, a negative result that we believe is valuable for guiding future research.